# Adaptive Debiased SGD in High-dimensional Generalized Linear Models with Streaming Data

### v2

@@ Abstract
Online statistical inference facilitates real-time analysis of sequentially collected data, making it different from traditional methods that rely on static datasets. This paper introduces a novel approach to online inference in high-dimensional generalized linear models, where we update regression coefficient estimates and their standard errors upon each new data arrival. In contrast to existing methods that either require full dataset access or large-dimensional summary statistics storage, our method operates in a single-pass mode, significantly reducing both time and space complexity. The core of our methodological innovation lies in an adaptive stochastic gradient descent algorithm tailored for dynamic objective functions, coupled with a novel online debiasing procedure. This allows us to maintain low-dimensional summary statistics while effectively controlling optimization errors introduced by the dynamically changing loss functions. We demonstrate, both methodologically and theoretically, that our method, termed the Adaptive Debiased Lasso (ADL), not only mitigates the need for the bounded individual probability condition but also significantly improves numerical performance. Numerical experiments demonstrate that the proposed ADL method consistently exhibits robust performance across various covariance matrix structures.
@@ 1 Introduction
In online inference, one of the main reasons to store the information matrix is that this quantity is pivotal in statistical inference and one can use such information directly to construct confidence interval for the target parameter(s). Moreover, online algorithms that leverage this second-order information can provide more accurate direction and step size towards the minimum, learning to faster convergence in some scenarios (Toulis and Airold,, 2017). However, as mentioned earlier, storing a $p\times p$ information matrix is generally not preferred in high-dimensional settings. As an alternative approach, the stochastic gradient descent (SGD) algorithm (Robbins and Monro,, 1951) fits the online setting well due to its computational and memory efficiency. In particular, SGD only requires one pass over the data, and its space complexity can reach $\mathcal{O}(p)$. The pioneering work that uses SGD for statistical inference can be traced back to Polyak and Juditsky, (1992) and it has also been applied to the low-dimensional models thereafter (Toulis and Airold,, 2017; Fang,, 2019; Chen et al.,, 2024). Recently, Chen et al., (2020) incorporate the SGD into high-dimensional inference task. Specifically, they combine the regularization annealed epoch dual averaging (RADAR) algorithm, one variant of SGD (Agarwal et al.,, 2012), with debiasing techniques (Zhang and Zhang,, 2014; van de Geer et al.,, 2014) in the linear models. However, their method requires access to the entire dataset in the debiasing step. Lately, Han et al., (2024) propose a new online debiased SGD algorithm in high-dimensional linear models that is a one-pass algorithm with space complexity $\mathcal{O}(p)$. While this algorithm has many desirable properties, it is not directly applicable to high-dimensional nonlinear models such as generalized linear models. In summary, to the best of our knowledge, existing online inference methods for high-dimensional generalized linear models are either not one-pass or require a space complexity of $\mathcal{O}(p^{2})$.
@@
The goal of this paper is to develop an online statistical inference method for high-dimensional generalized linear models. In particular, the proposed method is able to sequentially update the regression coefficient estimates and their standard errors upon the arrival of a new data point, without retrieving historical raw data of space complexity $\mathcal{O}(np)$ or saving summary statistics of space complexity $\mathcal{O}(p^{2})$. It is worth mentioning that the generalized linear model is quite different from the linear model, thus the methods in Chen et al., (2020) and Han et al., (2024) are not easily extendable to generalized linear models. The major challenges come from two aspects. First, the objective function in the online inference framework dynamically changes as the algorithm proceeds forward. This is fundamentally different from the RADAR algorithm where the objective function is fixed. As a result, the change in the objective function introduces an additional error term which should be carefully controlled to establish asymptotic normality. Second, due to the nonlinear nature of the generalized linear models, it is nontrivial to develop a one-pass algorithm that utilizes summary statistics of low-space complexity. In both the first-order gradient and Hessian matrix of the loss function, parameter estimates and data are intertwined together, making a direct decomposition infeasible.
@@
To address these challenges, we propose a new online debiasing procedure with two key components. First, we propose an adaptive RADAR algorithm that allows the underlying objective function change dynamically. More importantly, the optimization errors induced from the dynamically changed objective functions is still controllable, which provides the basis to further conduct valid statistical inference. Second, to avoid reaccessing and recomputing with historical raw data, we propose to linearize the nonlinear form of the first-order gradient via Taylor approximation. Such a linearization step is highly non-trivial. It is clear that second-order approximation can effectively reduce approximation errors, but it will involve Hessian matrix with space complexity of $\mathcal{O}(p^{2})$. Our key idea is to construct the summary statistics by combining the output from the proposed adaptive RADAR algorithm and the second-order information to reduce the dimensionality. In particular, these summary statistics require only $\mathcal{O}(p)$ space complexity but still retain the key information to conduct inference.
@@
Our contributions can be summarized from both methodological and theoretical aspects. Methodologically, we propose an online inference method for high-dimensional inference, named the adaptive debiased lasso (ADL). The proposed method offers two main computational benefits. First, it is designed to operate in a single-pass mode, resulting in low time complexity. Second, this method does not require storing the raw data or large summary statistics, and its space complexity is only of $\mathcal{O}(p)$, which is comparable to the complexity of storing a single data point. Thus, our method can handle ultra-high dimensional data effectively. In our real data analysis, we demonstrate that our method can handle datasets with millions of covariates on a standard personal computer. By contrast, alternative methods requiring access to the entire raw dataset or large summary statistics fail in similar scenarios due to memory constraints. A detailed comparison between our method and relevant existing literature is presented in Table 1.
In our theoretical work, we have removed the bounded individual probability condition – a condition commonly imposed in the literature to establish the asymptotic theorem in high-dimensional settings. This condition typically applies constraints to the sample data. Instead, our approach shifts focus from statistical error to optimization error, especially errors arising from the SGD algorithm. Notably, the objective function for SGD is based on the loss function at the population level rather than the sample level, so that we can avoid making assumptions on the sample data. We believe this SGD approach offers a novel perspective to handle many other statistical problems that are of independent interest. In addition, removing such a technical condition is not only a theoretical advancement but also improves the numerical performance. As demonstrated in Cai et al., (2023), the bounded individual probability condition is easily violated when the variance of the covariate is large, deteriorating the performance of most conventional high-dimensional inference methods. In contrast, our method consistently performs well even with varying covariance matrix structures, demonstrating robustness under different scenarios.
@@
Notation: We use $\vee$ and $\wedge$ to denote the $max$ and $min$ operator, respectively. For a random variable $Z$, its sub-Gaussian norm is defined as $\lVert Z\lVert_{\psi_{2}}=\sup_{q\geq 1}q^{-1/2}(\mathbb{E}|Z|^{q})^{1/q}$; For a random vector $\bm{Z}$, its sub-Gaussian norm is defined as $\lVert\bm{Z}\lVert_{\psi_{2}}=\sup_{\lVert\bm{v}\lVert_{2}=1}\lVert\bm{v}^{T}\bm{Z}\lVert_{\psi_{2}}$. We say $\bm{Z}$ is sub-Gaussian if its sub-Gaussian $\lVert\bm{Z}\lVert_{\psi_{2}}$ norm is finite. We also call a constant $c$ the universal constant if $c$ does not depend on $n$ and $p$. For $m\in\mathbb{N}$, we write $\{1,\cdots,m\}$ as $[m]$. We use $\mathcal{O}_{p}(\cdot)$ and $o_{p}(\cdot)$ to represent the equivalent and higher-order relation between two random sequences in probability, respectively. For a deterministic sequence, we use $\mathcal{O}(\cdot)$ in place of $\mathcal{O}_{p}(\cdot)$. We use $\lesssim$ and $\gtrsim$ to denote the asymptotic inequality relations.
@@ 2 Methodology
Suppose that $n$ data points $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ arrive sequentially and each of them is an independent and identically distributed copy of $(\bm{X},Y)$, which are sampled from a generalized linear model:
@@
Then the lasso estimator is defined as $\widetilde{\bm{\beta}}^{(n)}=\arg\min_{\bm{\beta}\in\mathbb{R}^{p}}F_{n}(\bm{\beta};\lambda_{n}).$ Since $\widetilde{\bm{\beta}}^{(n)}$ does not have a tractable limiting distribution (Zhao and Yu,, 2006), the debiased step was considered. Specifically, for some $j\in[p]$, the debiased estimator is calculated via
$\widetilde{\bm{\beta}}^{(n)}_{j,de}=\widetilde{\bm{\beta}}_{j}^{(n)}-\{\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}\}^{\intercal}f_{n}(\widetilde{\bm{\beta}}^{(n)}),$ (3)
where $f_{n}(\widetilde{\bm{\beta}}^{(n)})=\nabla F_{n}(\widetilde{\bm{\beta}}^{(n)})$ is the gradient of the negative log-likelihood function, and $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$, known as the the nodewise lasso estimator, serves as an estimator of the $j$-th column in the inverse of Hessian matrix $\nabla^{2}F_{n}(\widetilde{\bm{\beta}}^{(n)})$. Zhang and Zhang, (2014) and van de Geer et al., (2014) show that such a debiased estimator converges to the normal distribution under the linear models and generalized linear models respectively.
It is straightforward to check whether (3) could fit into the online learning framework. Apparently, the calculations of $\widetilde{\bm{\beta}}^{(n)}$, $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$, and $f_{n}(\widetilde{\bm{\beta}}^{(n)})$ all require re-accessing and recomputing with the full dataset. For this reason, we consider developing real online methods and algorithms that update the debiased lasso estimator without revisiting historical data. Based on existing literature in online lasso estimation, online alternative for estimating $\widetilde{\bm{\beta}}^{(n)}$ is readily available (Agarwal et al.,, 2012). Nonetheless, as indicated by (3), there are two remaining components we need to deal with to conduct a one-step estimation that fits into an online statistical inference framework.
In the following, we develop an online method to conduct statistical inference. We first introduce the RADAR algorithm used to conduct online lasso estimation in Section 2.1. Specifically, we develop an adaptive RADAR algorithm to find online alternatives to replace $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$ in Section 2.2. We then propose an approximated debiasing technique to replace $f_{n}(\widetilde{\bm{\beta}}^{(n)})$. The resulting estimator, named adaptive debiased lasso (ADL) estimator, as well as its standard error, are presented in Sections 2.32.4. For illustration, we present the entire procedure in Figure 1. The online lasso estimation algorithm is marked in blue. Our proposed method, including the adaptive RADAR and approximated debiasing technique, is highlighted in orange. We wish to emphasize that our procedure is one-pass and only requires $\mathcal{O}(p)$ space complexity.
@@ 2.1 Online lasso with RADAR
In this subsection, we introduce an online one-pass approach to finding the lasso estimator without retrieving or recomputing historical data. A natural choice that serves this purpose is the SGD (Robbins and Monro,, 1951). Although SGD can obtain an estimator of $\bm{\beta}^{*}$, its convergence rate is unsatisfactory in high-dimensional settings. For example, Han et al., (2024) show that interval estimation using the vanilla SGD may not have desirable performance such as attaining nominal coverage probability. To address this issue, we adopt the regularization annealed epoch dual averaging (RADAR) algorithm (Agarwal et al.,, 2012), a variant of the vanilla SGD.
Different from the vanilla SGD, the RADAR algorithm incorporates the multi-epoch technique. The multi-epoch technique splits a sequence of $n$ data points into $K$ epochs according to a sequence of epoch length schedule $\{T_{k}\}_{k\in[K]}$, where $k$-th epoch contains data of sample size $T_{k}$ and $\sum_{k=1}^{K}T_{k}=n$. Let $n_{k}=\sum_{i=1}^{k}T_{i}$ denote the cumulative sample size up to the $k$-th epoch. In the $k$-th epoch, the RADAR algorithm runs $T_{k}$ rounds of the proximal SGD with dual averaging (Xiao,, 2009) and outputs the estimate $\widehat{\bm{\beta}}^{(n_{k})}$ at the end of the epoch. The output $\widehat{\bm{\beta}}^{(n_{k})}$ is then used as an initial value in the $(k+1)$-th epoch. The multi-epoch method accelerates the convergence rate by dropping the information received in the previous epochs. Since the RADAR algorithm outputs the estimate only at the end of each epoch, $\widehat{\bm{\beta}}^{(i)}$ is well-defined only for $i=n_{k},k\in[K]$. For convenience, we define the estimates within the $(k+1)$-th epoch as
@@
It is worth noting that $\widehat{{\bm{\beta}}}^{(n)}$, the final output from the RADAR algorithm, is no longer the minimizer of $F_{n}(\bm{\beta};\lambda_{n})$ defined in (2). In fact, without accessing the entire dataset, it is hard to achieve the optimality as the offline estimator. Recall $\widetilde{\bm{\beta}}^{(n)}$ is the minimizer of $F_{n}(\bm{\beta};\lambda_{n})$. Instead of the statistical error $\lVert\widetilde{\bm{\beta}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$, we need to consider the optimization error from the RADAR algorithm, that is $\lVert\widehat{{\bm{\beta}}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$. Under some regularity conditions, we can conclude that the optimization error $\lVert\widehat{{\bm{\beta}}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$ achieves the same rate as the statistical error $\lVert\widetilde{\bm{\beta}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$. A detailed description of the RADAR algorithm is included in the supplementary material.

### v3

# Adaptive Debiased Lasso in High-dimensional Generalized Linear Models with Streaming Data

@@ Abstract
Online statistical inference facilitates real-time analysis of sequentially collected data, making it different from traditional methods that rely on static datasets. This paper introduces a novel approach to online inference in high-dimensional generalized linear models, where we update regression coefficient estimates and their standard errors upon each new data arrival. In contrast to existing methods that either require full dataset access or large-dimensional summary statistics storage, our method operates in a single-pass mode, significantly reducing both time and space complexity. The core of our methodological innovation lies in an adaptive stochastic gradient descent algorithm tailored for dynamic objective functions, coupled with a novel online debiasing procedure. This allows us to maintain low-dimensional summary statistics while effectively controlling the optimization error introduced by the dynamically changing loss functions. We establish the asymptotic normality of our proposed Adaptive Debiased Lasso (ADL) estimator. We conduct extensive simulation experiments to show the statistical validity and computational efficiency of our ADL estimator across various settings. Its computational efficiency is further demonstrated via a real data application to the spam email classification.
@@ 1 Introduction
In online inference, one of the main reasons to store the information matrix is that this quantity is pivotal in statistical inference and one can use such information directly to construct confidence interval for the target parameter(s). Moreover, online algorithms that leverage this second-order information can provide more accurate direction and step size towards the minimum, leading to faster convergence in some scenarios (Toulis and Airold,, 2017). However, as mentioned earlier, storing a $p\times p$ information matrix is generally not preferred in high-dimensional settings. As an alternative approach, the stochastic gradient descent (SGD) algorithm (Robbins and Monro,, 1951) fits the online setting well due to its computational and memory efficiency. In particular, SGD only requires one pass over the data, and its space complexity can reach $\mathcal{O}(p)$. The pioneering work that uses SGD for statistical inference can be traced back to Polyak and Juditsky, (1992) and this idea has also been applied to the low-dimensional models thereafter (Toulis and Airold,, 2017; Fang,, 2019; Chen et al.,, 2024). Recently, Chen et al., (2020) incorporate the SGD into high-dimensional inference tasks. Specifically, they combine the regularization annealed epoch dual averaging (RADAR) algorithm, one variant of SGD (Agarwal et al.,, 2012), with debiasing techniques (Zhang and Zhang,, 2014; van de Geer et al.,, 2014) in the linear models. However, their method requires access to the entire dataset in the debiasing step. Lately, Han et al., (2024) propose a new online debiased SGD algorithm in high-dimensional linear models that is a one-pass algorithm with space complexity $\mathcal{O}(p)$. While this algorithm has many desirable properties, it is not directly applicable to high-dimensional nonlinear models such as generalized linear models. In summary, to the best of our knowledge, existing online inference methods for high-dimensional generalized linear models are either not one-pass or require a space complexity of $\mathcal{O}(p^{2})$.
@@
The goal of this paper is to develop an online statistical inference method for high-dimensional generalized linear models. In particular, the proposed method is able to sequentially update the regression coefficient estimates and their standard errors upon the arrival of a new data point, without retrieving historical raw data of space complexity $\mathcal{O}(np)$ or saving summary statistics of space complexity $\mathcal{O}(p^{2})$. It is worth mentioning that the generalized linear models are quite different from the linear models, thus the methods in Chen et al., (2020) and Han et al., (2024) are not easily extendable to generalized linear models. The major challenges come from two aspects. First, the objective function in the online inference framework dynamically changes as the algorithm proceeds forward. This is fundamentally different from the RADAR algorithm where the objective function is fixed. As a result, the change in the objective function introduces an additional error term which should be carefully controlled to establish asymptotic normality. Second, due to the nonlinear nature of the generalized linear models, it is nontrivial to develop a one-pass algorithm that utilizes summary statistics of low-space complexity. In both the first-order gradient and the Hessian matrix of the loss function, parameter estimates and data are intertwined together, making a direct decomposition infeasible.
@@
To address these challenges, we propose a new online debiasing procedure with two key components. First, we propose an adaptive RADAR algorithm that allows the underlying objective function to change dynamically. More importantly, the optimization error induced from the dynamically changed objective functions is still controllable, which provides the basis to further conduct valid statistical inference. Second, to avoid reaccessing and recomputing with historical raw data, we propose to linearize the nonlinear form of the first-order gradient via Taylor approximation. Such a linearization step is highly non-trivial. It is clear that second-order approximation can effectively reduce approximation error, but it will involve a Hessian matrix with space complexity of $\mathcal{O}(p^{2})$. Our key idea is to construct the summary statistics by combining the output from the proposed adaptive RADAR algorithm and the second-order information to reduce the dimensionality. In particular, these summary statistics require only $\mathcal{O}(p)$ space complexity but still retain the key information to conduct inference.
@@
Our contributions can be summarized from both methodological and theoretical aspects. Methodologically, we propose an online inference method for high-dimensional inference, named the adaptive debiased lasso (ADL). The proposed method offers two main computational benefits. First, it is designed to operate in a single-pass mode, resulting in low time complexity. Second, this method does not require storing the raw data or large summary statistics, and its space complexity is only $\mathcal{O}(p)$, which is comparable to the complexity of storing a single data point. Thus, our method can handle ultra-high dimensional data effectively. In our real data example, we demonstrate that our method can handle datasets with millions of covariates on a standard personal computer. By contrast, alternative methods requiring access to the entire raw dataset or large summary statistics fail in similar scenarios due to memory constraints. A detailed comparison between our method and relevant existing literature is presented in Table 1. Theoretically, we establish the asymptotic normality of the proposed ADL estimator, which is an output from an online one-pass algorithm for high-dimensional generalized linear models. The asymptotic normality is built upon the consistency of two estimators, one is the lasso estimator from the RADAR algorithm and the other is the nodewise lasso estimator from the adaptive RADAR algorithm. The latter requires careful analysis of the cumulative error in a stochastic optimization procedure with a dynamically changing objective function. Moreover, unlike the conventional high-dimensional inference in offline settings, the online debiasing step involves a Taylor approximation step, and we show that the cumulative residual as a result of this approximation is negligible.
@@
Notation: We use $\vee$ and $\wedge$ to denote the $max$ and $min$ operator, respectively. For a random variable $Z$, its sub-Gaussian norm is defined as $\lVert Z\lVert_{\psi_{2}}=\sup_{q\geq 1}q^{-1/2}(\mathbb{E}|Z|^{q})^{1/q}$; For a random vector $\bm{Z}$, its sub-Gaussian norm is defined as $\lVert\bm{Z}\lVert_{\psi_{2}}=\sup_{\lVert\bm{v}\lVert_{2}=1}\lVert\bm{v}^{T}\bm{Z}\lVert_{\psi_{2}}$. We say $\bm{Z}$ is sub-Gaussian if its sub-Gaussian norm $\lVert\bm{Z}\lVert_{\psi_{2}}$ is finite. We also call a constant $c$ the universal constant if $c$ does not depend on $n$ and $p$. For $m\in\mathbb{N}$, we write $\{1,\cdots,m\}$ as $[m]$. For deterministic sequences, we use $\mathcal{O}(\cdot)$ and $o(\cdot)$ to represent the Bachmann–Landau asymptotic notation. For random sequences, we use $\mathcal{O}_{\mathbb{P}}(\cdot)$ and $o_{\mathbb{P}}(\cdot)$ in place of $\mathcal{O}(\cdot)$ and $o(\cdot)$ respectively. We use $\lesssim$ and $\gtrsim$ to denote the asymptotic inequality relations. For a set $S$, we use $S^{c}$ as its complement. For a square matrix $\bm{\Sigma}$, we denote its minimum and maximum eigenvalues by $\Lambda_{\min}(\bm{\Sigma})$ and $\Lambda_{\max}(\bm{\Sigma})$ respectively.
@@ 2 Methodology
Suppose that $n$ data points $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ arrive sequentially and each of them is an independent and identically distributed copy of $(\bm{X},Y)$, sampled from a generalized linear model:
@@
Then the lasso estimator is defined as $\widetilde{\bm{\beta}}^{(n)}=\arg\min_{\bm{\beta}\in\mathbb{R}^{p}}F_{n}(\bm{\beta};\lambda_{n}).$ Since $\widetilde{\bm{\beta}}^{(n)}$ does not have a tractable limiting distribution (Zhao and Yu,, 2006), the debiased step is considered. Specifically, for some $j\in[p]$, the debiased estimator is calculated via
$\widetilde{\bm{\beta}}^{(n)}_{j,de}=\widetilde{\bm{\beta}}_{j}^{(n)}-\frac{\{\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}\}^{\intercal}\{\nabla F_{n}(\widetilde{\bm{\beta}}^{(n)})\}}{~{}~{}~{}\{\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}\}^{\intercal}\{\nabla^{2}F_{n}(\widetilde{\bm{\beta}}^{(n)})\}_{\cdot,j}},$ (3)
where $\nabla F_{n}(\widetilde{\bm{\beta}}^{(n)})$ is the gradient of the negative log-likelihood function, $\{\nabla^{2}F_{n}(\widetilde{\bm{\beta}}^{(n)})\}_{\cdot,j}$ is the $j$th column of the Hessian matrix, and $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$ is known as the the nodewise lasso estimator. Zhang and Zhang, (2014) and van de Geer et al., (2014) show that such a debiased estimator converges to the normal distribution under the linear models and generalized linear models respectively.
It is straightforward to check whether (3) could fit into the online learning framework. Apparently, the calculations of $\widetilde{\bm{\beta}}^{(n)}$, $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$, $\nabla F_{n}(\widetilde{\bm{\beta}}^{(n)})$, and $\{\nabla^{2}F_{n}(\widetilde{\bm{\beta}}^{(n)})\}_{\cdot,j}$ all require re-accessing and recomputing with the full dataset. For this reason, we consider developing real online methods and algorithms that update the debiased lasso estimator without revisiting historical data. Based on existing literature in online lasso estimation, an online alternative for $\widetilde{\bm{\beta}}^{(n)}$ is readily available (Agarwal et al.,, 2012). Nonetheless, as indicated by (3), there are three remaining components we need to deal with to conduct a one-step estimation that fits into an online statistical inference framework. In the following, we develop an online method to conduct statistical inference. We first introduce the RADAR algorithm used to conduct online lasso estimation in Section 2.1. Then, we develop an adaptive RADAR algorithm to find an online alternative to replace $\widetilde{\bm{\gamma}}^{(n)}_{\cdot,j}$ in Section 2.2. Furthermore, we propose an approximated debiasing technique to replace $\nabla F_{n}(\widetilde{\bm{\beta}}^{(n)})$ and $\{\nabla^{2}F_{n}(\widetilde{\bm{\beta}}^{(n)})\}_{\cdot,j}$. The resulting estimator, named adaptive debiased lasso (ADL) estimator, as well as its standard error, are presented in Sections 2.32.4. For illustration, we present the entire procedure in Figure 1. The online lasso estimation algorithm is marked in blue. Our proposed method, including the adaptive RADAR and approximated debiasing technique, is highlighted in orange. We wish to emphasize that our procedure is one-pass and only requires $\mathcal{O}(p)$ space complexity.
@@ 2.1 Online lasso with RADAR
In this section, we introduce an online one-pass approach to finding the lasso estimator without retrieving or recomputing historical data. A natural choice that serves this purpose is the SGD (Robbins and Monro,, 1951). Although SGD can obtain an estimator of $\bm{\beta}^{*}$, its convergence rate is unsatisfactory in high-dimensional settings. Specifically, Shalev-Shwartz and Tewari, (2011) show that the convergence rate of SGD in high-dimensional settings with $\ell_{1}$-norm penalty is $\mathcal{O}(n^{-1/4})$, which is slower than $\mathcal{O}(n^{-1/2})$. Thus, interval estimation using the vanilla SGD may not have desired theoretical guarantee. In addition,  Han et al., (2024) also show numerically that interval estimators with the vanilla SGD may not attain nominal coverage probability. To address this issue, we adopt the regularization annealed epoch dual averaging (RADAR) algorithm (Agarwal et al.,, 2012), a variant of the vanilla SGD.
Different from the vanilla SGD, the RADAR algorithm incorporates the multi-epoch technique. The multi-epoch technique splits a sequence of $n$ data points into $K$ epochs according to a sequence of epoch length schedule $\{T_{k}\}_{k\in[K]}$, where the $k$th epoch contains data of sample size $T_{k}$ and $\sum_{k=1}^{K}T_{k}=n$. Let $n_{k}=\sum_{i=1}^{k}T_{i}$ denote the cumulative sample size up to the $k$th epoch. In the $k$th epoch, the RADAR algorithm runs $T_{k}$ rounds of the proximal SGD with dual averaging (Xiao,, 2009) and outputs the estimate $\widehat{\bm{\beta}}^{(n_{k})}$ at the end of the epoch. The output $\widehat{\bm{\beta}}^{(n_{k})}$ is then used as an initial value in the $(k+1)$th epoch. This multi-epoch method accelerates the convergence rate by dropping the information received in the previous epochs. Since the RADAR algorithm outputs the estimate only at the end of each epoch, $\widehat{\bm{\beta}}^{(i)}$ is well-defined only for $i=n_{k},k\in[K]$. For convenience, we define the estimates within the $(k+1)$th epoch as
@@
It is worth noting that $\widehat{{\bm{\beta}}}^{(n)}$, the final output from the RADAR algorithm, is no longer the minimizer of $F_{n}(\bm{\beta};\lambda_{n})$ defined in (2). In fact, without accessing the entire dataset, it is hard to achieve the optimality as the offline estimator. Recall $\widetilde{\bm{\beta}}^{(n)}$ is the minimizer of $F_{n}(\bm{\beta};\lambda_{n})$. Instead of the statistical error $\lVert\widetilde{\bm{\beta}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$, we need to consider the optimization error from the RADAR algorithm, that is $\lVert\widehat{{\bm{\beta}}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$. Under some regularity conditions, we can conclude that the optimization error $\lVert\widehat{{\bm{\beta}}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$ achieves the same rate as the statistical error $\lVert\widetilde{\bm{\beta}}^{(n)}-\bm{\beta}^{*}\lVert_{1}$ in terms of the sample size $n$. A detailed description of the RADAR algorithm is included in the supplementary material.